A projectile is fired at an angle αwith initial velocity v0. If it experiences the 

resistive force 𝐹𝑟 = −𝑘𝑣

2

, find the expression for distance covered x as a 

function of vx ant time. Also discuss the special cases.


1
Expert's answer
2021-02-15T07:28:10-0500

Answer

For projectile motion

"u_x=v_0cos\alpha"

"u_y=v_0sin\alpha"

Force in x direction

"F_x=-kv_x^2=ma_x"

So acceleration

"a_x=\frac{dv_x}{dt}=-\frac{ kv_x^2}{m}"


So using integration limits

"\int_{v_0cos\alpha}^{v_x}\frac{dv_x}{v_x^2}=\int _0^t\frac{-kdt}{m}"

So we get


"[\frac{1}{v_x}]_{v_0cos\alpha}^{v_0}=-\frac{kt}{m}"

So in terms of vx and t

"v_x=-\frac{m}{kt}\\v_x\propto\frac{1}{t}"